# Extremum Seeking Tracking for Derivative-free Distributed Optimization

### v4

@@ IV-B Averaged System Analysis
There exist $P=P^{\top}\in{\mathbb{R}}^{n_{\xi}\times n_{\xi}}$, $a_{1},a_{2},c_{1}>0$, and $\gamma_{0}>0$ such that, for any $\gamma\in(0,\gamma_{0})$, along the trajectories of (28) it holds
@@
for all $\xi_{\text{avg}}\in{\mathbb{R}}^{n_{\xi}}$. $\square$
@@
Notice that Lemma 2 proves the algorithm (8) linearly converges to the minimizer of (1), since (28) is an equivalent formulation of (8). $\square$
With this result at hand, we analyze the impact of $u(\cdot)$ thus obtaining the stability properties for the averaged system (27).
@@
(i) $\xi_{\text{avg}}^{t}\in\mathcal{B}_{\sqrt{a_{2}/a_{1}}r_{\xi}}$ for all $t\geq 0$,
@@ IV-C Proof of Theorem 1
Despite averaging tools for discrete-time systems are already present in the literature, see [39] for example, we got the inspiration from continuous-time averaging [40, Ch. 10] and [41] for elaborating the proof of Theorem 1 to make clear how $\gamma$ affects the closeness of the trajectories of (18) and (27). Since Assumptions 12, and 3 hold, we apply Lemma 2 to claim that there exists $P=P^{\top}\in{\mathbb{R}}^{n_{\xi}\times n_{\xi}}$, $\gamma_{0}>0$, and $a_{1},a_{2},c_{1}>0$ such that, if $\gamma\in(0,\gamma_{0})$, the conditions (29) are satisfied. Then, we evaluate the norm of the initial conditions of system (18) and (27), i.e., $\left\|\xi^{0}\right\|=\left\|\xi^{0}_{\text{avg}}\right\|$. By using the definition of $\xi$ (cf. (15)), the changes of variables (12) and (14), and the triangle inequality, we get
@@
Once the initial distance from the origin has been evaluated, we choose any $\bar{\rho}>0$, set $c_{2}:=\sqrt{a_{2}/a_{1}}$, and choose any $\epsilon\in(0,\bar{\rho}(1+c_{2})/2)$. Then, we pick $\rho\in(0,(\bar{\rho}/2-(1+c_{2})\epsilon)/c_{2})$, $c_{1}^{\prime}\in(0,c_{1})$, and use the matrix $P$ satisfying (29) to apply Lemma 3. Specifically, we claim that there exist $c_{3}>0$, $\delta^{\star}_{1}\in(0,1]$, and $\gamma_{0}>0$ such that, for any $\delta\in(0,\delta^{\star}_{1})$ and $\gamma\in(0,\min\{\gamma_{0},\}]$, it holds $\xi_{\text{avg}}^{t}\in\mathcal{B}_{c_{2}r_{\xi}}$ for all $t\geq 0$ and
@@
for all $\xi_{\text{avg}}^{t}$ such that $\left\|\xi_{\text{avg}}^{t}\right\|\geq\rho$. Now, we proceed by finding a bound $\gamma_{1}>0$ such that, for any $\gamma\in(0,\gamma_{1})$, we guarantee the $\epsilon$-closeness between the state $\xi_{\text{avg}}^{t}$ of the averaged system (20) and the one of the original system (18), i.e., that $\|\xi^{t}-\xi_{\text{avg}}^{t}\|\leq\epsilon$ holds true for all $t\geq 0$. To this end, let us introduce
@@
Then, let $r_{\xi}^{\prime}:=c_{2}r_{\xi}$ and define $\Delta:=\delta\sqrt{Nn}$. Under the assumption of $\xi^{t}\in\mathcal{B}_{r_{\xi}^{\prime}+\epsilon}$ for all $t\geq 0$ (later verified by a proper selection of $\gamma$), we claim that the arguments of the functions $f_{i}$ and their derivatives (embedded into the definitions of $\phi(t,\cdot)$ and $\phi_{\text{avg}}(\cdot)$ and their derivatives) lie into the compact set $\mathcal{B}_{r^{\prime}_{\xi}+\epsilon+\Delta}$. Thus, since the functions $f_{i}$ and its derivatives are continuous (cf. Assumption 3) and the functions $\phi(\cdot,\cdot)$ and $\nu(\cdot,\cdot)$ are periodic in the first argument, we define
@@
Consequently, for all $\xi,\xi^{\prime}\in\mathcal{B}_{r_{\xi}^{\prime}+\epsilon}$ and $t\geq 0$, it holds
@@
Let $\gamma_{2}:=\tfrac{\epsilon/3}{L_{\phi}^{2}2\left(1+2\tau_{\text{per}}\right)\theta^{\star}}$, $\gamma_{3}:=\tfrac{\epsilon/3}{2L_{\phi}(1+2\tau_{\text{per}})\exp(L_{\phi}\theta^{\star})}$,$\gamma_{4}:=\tfrac{\epsilon/3}{2L_{\phi}\tau_{\text{per}}}$, $\gamma_{1}:=\min\{\gamma_{0},\gamma_{2},\gamma_{3},\gamma_{4},1\}$. Pick $\gamma\in(0,\gamma_{1})$ such that $t^{\star}:=\frac{\theta^{\star}}{\gamma}\in{\mathbb{N}}$. This can be done without loss of generality since $\theta^{\star}$ is a design parameter. Then, we bound (38) as
$\|\xi^{t}-\xi_{\text{avg}}^{t}\|\leq\epsilon,$ (40)
for all $t\in[0,t^{\star}]$. As a consequence, since $\xi_{\text{avg}}^{t}\in\mathcal{B}_{r_{\xi}^{\prime}}$ for all $t\geq 0$, it holds ${\xi}^{t}\in\mathcal{B}_{r_{\xi}^{\prime}+\epsilon}$ for all $t\in[0,t^{\star}]$, i.e., we have verified that the bounds (34) can be used into the interval $[0,t^{\star}]$. Moreover, the exponential law (31) and the expression of $\theta^{\star}$ (cf. (39)) ensure that it holds
@@
where in $(a)$ we combined (40) and (41). The inequality (42) guarantees that $\xi^{t^{\star}}\in\mathcal{B}_{\frac{\bar{\rho}}{2c_{2}}}$, hence we proved that the trajectories of (27) enters into $\mathcal{B}_{\frac{\bar{\rho}}{2c_{2}}}$ with linear rate. Next, in order to show that $\xi^{t}\in\mathcal{B}_{\bar{\rho}/2}$ for all $t\geq t^{\star}$, we divide the set of natural numbers in intervals as $\mathbb{N}=[0,\,t^{\star}]\cup[t^{\star},2t^{\star}]\cup\dots$. Define $\psi_{\text{avg}}(k+t^{\star},\xi^{t^{\star}})$ as the solution to (27) for $\xi_{\text{avg}}^{0}=\xi^{t^{\star}}$ and $k\in[0,t^{\star}]$. Thus, at the beginning of each interval $[t^{\star},2t^{\star}]$, the initial condition of (27) coincides with the one of $\psi_{\text{avg}}(k+t^{\star},\xi^{t^{\star}})$ and lies into $\mathcal{B}_{\bar{\rho}/2}\subseteq\mathcal{B}_{r_{\xi}}$. Thus, we apply the same arguments above to guarantee that
$\|\xi^{k+t^{\star}}-\psi_{\text{avg}}(k+t^{\star},\xi^{t^{\star}})\|\leq\epsilon,$
for all $\gamma\in(0,\gamma^{\star})$, $k\in[0,t^{\star}]$, and $\psi_{\text{avg}}(2t^{\star},\xi^{t^{\star}})\in\mathcal{B}_{(\bar{\rho}/2-\epsilon)/c_{2}}$. By using Lemma 3, we guarantee that system (27) cannot escape from $\mathcal{B}_{\bar{\rho}/2-\epsilon}$, namely $\xi_{\text{avg}}^{t}\in\mathcal{B}_{\bar{\rho}/2-\epsilon}$ for all $t\geq t^{\star}$. Thus, we get $\xi^{t}\in\mathcal{B}_{\bar{\rho}/2}$ for all $t\in[t^{\star},2t^{\star}]$. By recursively applying the same arguments above for each interval $[jt^{\star},(j+1)t^{\star}]$ with $j=2,3,\dots$ and using $\left\|x_{i}^{t}-x^{\star}\right\|\leq\left\|\xi^{t}\right\|$ for all $i\in\{1,\ldots,N\}$ and $t\geq 0$, we get
@@
for all $i\in\{1,\ldots,N\}$ and $t\geq 0$. The change of coordinates (6), (43), and the triangle inequality lead to
$\left\|w_{i}^{t}-x^{\star}\right\|$ $\leq\frac{\bar{\rho}}{2}+\delta\left\|d_{i}^{t}\right\|\stackrel{{\scriptstyle(a)}}{{\leq}}\bar{\rho}/2+\delta\sqrt{n},$ (44)
@@ V Numerical Computations on Distributed personalized optimization
In several engineering applications, a problem of interest consists of optimizing a performance metric while keeping into account user discomfort terms [9, 44]. In these scenarios, the user discomfort term is usually not known in advance but can be only accessed by measurements. Specifically, we associate to each agent $i\in\{1,\ldots,N\}$ a cost function in the form $f_{i}(w)=w^{\top}Q_{i}w+r_{i}^{\top}w+\log(\sum_{\ell=1}^{n}a_{i\ell}e^{b_{i\ell}w_{\ell}})$ with $Q_{i}=Q_{i}^{\top}\in^{n\times n}$, $r_{i}\in^{n}$ and $a_{i\ell},b_{i\ell}$, for all $\ell\in\{1,\ldots,n\}$. In the following, we provide different sets of simulations to study in detail the features of Extremum Seeking Tracking. In particular, each set consists of Monte Carlo simulations over $20$ randomly generated scenarios in which each one differs from the other in terms of cost functions, communication graphs, and algorithmic variables’ initialization. In particular, unless differently stated, in each trial the agents communicate according to Erdős-Rényi random graphs (see, e.g., [45]) with edge probabilities equal to $0.2$. As for the problem parameters, for each trial and all $i\in\{1,\ldots,N\}$, we generate each matrix $Q_{i}$ by pre- and post-multiplying a diagonal matrix (whose diagonal elements are randomly extracted from the interval $[10^{-3},5\cdot 10^{-3}]$ with uniform probability) with an orthonormal matrix (randomly generated by extracting its elements from the interval $[0,1]$ with uniform probability) and its transpose, respectively. Further, for all $i\in\{1,\ldots,N\}$ and $\ell\in\{1,\ldots,n\}$, we randomly extract the components of $r_{i}$ within the interval $[-10^{-2},3\cdot 10^{-2}]$ and the parameters $a_{i\ell},b_{i\ell}$ within the interval $[0,10^{-3}]$ with a uniform probability. In each set of simulations, we choose the parameters $\tau_{i_{p}}$ and $\phi_{i_{p}}$ according to the following procedure. For all $i\in\{1,\ldots,N\}$, we take $\phi_{i_{p}}=\dfrac{\pi}{4}\left(1+(-1)^{p}\right)$ for all $p=1,\dots,n$, while $\tau_{i_{p}}$ have been chosen as the first $\lfloor(n+1)/2\rfloor$ elements of odd numbers greater than $3$ since it is possible to show that such a set of frequencies satisfy (3). Roughly speaking, $\phi_{i,p}=0$ and $\phi_{i,p}=\pi/2$ are used to create orthogonal functions with the same frequency. It is important to note that this selection is not unique. As for the algorithm parameters $\gamma$ and $\delta$, due to the complexity of the dependencies of the bounds $\gamma^{\star}$ and $\delta^{\star}$ provided in Theorem 1, we choose them via a trial-and-error procedure. In each set of simulations, the performance is evaluated by providing graphical results involving the relative errors ${|\sum_{i}f_{i}(\bar{x}^{t})-f(x^{\star})|}/{|f(x^{\star})|}$ and ${\|\bar{x}^{t}-x^{\star}\|}/{\|x^{\star}\|}$ achieved along the trials of the Monte Carlo simulations, where $\bar{x}^{t}:=\frac{1}{N}\sum_{i=1}^{N}x_{i}^{t}$. Simulations are performed using DISROPT [46], a Python package based on MPI to encode and simulate distributed optimization algorithms.
@@ V-A Monte Carlo simulations for large-scale problems
First, we perform numerical simulations over large-scale problems with networks made of $N=250$ agents. We consider different optimization variable sizes, namely $n=10,20$. Part of these performances has been run on the Marconi100 HPC Cluster of the Italian Cineca. We used $10$ nodes of the cluster and, for each node, we used $25$ cores and $4$ GPUs. The code has been adapted in order to perform part of the computation directly on GPUs. The results are shown in Fig. 2. In detail, increasing the number of agents increases the Lipschitz constant of the system to be averaged. Moreover, a larger domain of initial conditions also implies a potentially larger $L_{\phi}$ constant (cf. (33)). This implies smaller $\gamma^{\star}$, which, fixed the other parameters, makes the convergence slower. The decision variable dimension instead impacts the selection of the dither signal. A larger number of states implies a larger number of frequencies. This, in turn, means a longer time to estimate the gradient (cf. Lemma 1). Notice that, however, the accuracy of the final estimate is guaranteed by design. Indeed, since $\delta$ and $\gamma$ are designed on $\bar{\rho}$, the trajectories of (5) converge to a ball of radius $\bar{\rho}$ independently of the problem size.

### v6

@@ IV-B Averaged System Analysis
There exist $P=P^{\top}\in{\mathbb{R}}^{n_{\xi}\times n_{\xi}}$ and $a_{1},a_{2},c_{1},\gamma_{0}>0$ such that, for any $\gamma\in(0,\gamma_{0})$, along the trajectories of (28) it holds
@@
for all $\xi_{\text{avg}}^{t}\in{\mathbb{R}}^{n_{\xi}}$. $\square$
@@
Notice that Lemma 2 proves that algorithm (8) linearly converges to the minimizer of (1), since (28) is an equivalent formulation of (8). $\square$
With this result at hand, we analyze the impact of $u(\cdot)$ thus obtaining the stability properties of the averaged system (27).
@@
(i) $\xi_{\text{avg}}^{t}\in\mathcal{B}_{\sqrt{a_{2}/a_{1}}r_{\xi}}$ for all $t\in\mathbb{N}$,
@@ IV-C Proof of Theorem 1
Despite averaging tools for discrete-time systems are already present in the literature, see [39] for example, we got the inspiration from continuous-time averaging [40, Ch. 10] and [41] for elaborating the proof of Theorem 1 to make clear how $\gamma$ affects the closeness of the trajectories of (18) and (27). Since Assumptions 12, and 3 hold, we apply Lemma 2 to claim that there exist $P=P^{\top}\in{\mathbb{R}}^{n_{\xi}\times n_{\xi}}$ and $a_{1},a_{2},c_{1},\gamma_{0}>0$ such that, if $\gamma\in(0,\gamma_{0})$, the conditions (29) are satisfied. Then, we evaluate the norm of the initial conditions of system (18) and (27), i.e., $\left\|\xi^{0}\right\|=\left\|\xi^{0}_{\text{avg}}\right\|$. By using the definition of $\xi$ (cf. (15)), the changes of variables (12) and (14), and the triangle inequality, we get
@@
Once the initial distance from the origin has been evaluated, we choose any $\bar{\rho}>0$, set $c_{2}:=\sqrt{a_{2}/a_{1}}$, and choose any $\epsilon\in(0,\bar{\rho}(1+c_{2})/2)$. Then, we pick $\rho\in(0,(\bar{\rho}/2-(1+c_{2})\epsilon)/c_{2})$, $c_{1}^{\prime}\in(0,c_{1})$, and use the matrix $P$ satisfying (29) to apply Lemma 3. Specifically, we claim that there exist $c_{3}>0$, $\delta^{\star}_{1}\in(0,1]$, and $\gamma_{0}>0$ such that, for any $\delta\in(0,\delta^{\star}_{1})$ and $\gamma\in(0,\min\{\gamma_{0},1\})$, it holds $\xi_{\text{avg}}^{t}\in\mathcal{B}_{c_{2}r_{\xi}}$ for all $t\in\mathbb{N}$ and
@@
for all $\xi_{\text{avg}}^{t}$ such that $\left\|\xi_{\text{avg}}^{t}\right\|\geq\rho$. Now, we proceed by finding a bound $\gamma_{1}>0$ such that, for any $\gamma\in(0,\gamma_{1})$, we guarantee the $\epsilon$-closeness between the state $\xi_{\text{avg}}^{t}$ of the averaged system (20) and the one of the original system (18), i.e., that $\|\xi^{t}-\xi_{\text{avg}}^{t}\|\leq\epsilon$ holds true for all $t\in\mathbb{N}$. To this end, let us introduce
@@
Then, let $r_{\xi}^{\prime}:=c_{2}r_{\xi}$ and define $\Delta:=\delta\sqrt{Nn}$. Under the assumption of $\xi^{t}\in\mathcal{B}_{r_{\xi}^{\prime}+\epsilon}$ for all $t\in\mathbb{N}$ (later verified by a proper selection of $\gamma$), we claim that the arguments of the functions $f_{i}$ and their derivatives (embedded into the definitions of $\phi(t,\cdot)$ and $\phi_{\text{avg}}(\cdot)$ and their derivatives) lie into the compact set $\mathcal{B}_{r^{\prime}_{\xi}+\epsilon+\Delta}$. Thus, since the functions $f_{i}$ and its derivatives are continuous (cf. Assumption 3) and the functions $\phi(\cdot,\cdot)$ and $\nu(\cdot,\cdot)$ are periodic in the first argument, we define
@@
Consequently, for all $\xi,\xi^{\prime}\in\mathcal{B}_{r_{\xi}^{\prime}+\epsilon}$ and $t\in\mathbb{N}$, it holds
@@
Let $\gamma_{2}:=\tfrac{\epsilon/(3c_{2})}{L_{\phi}^{2}2\left(1+2\tau_{\text{per}}\right)\theta^{\star}}$, $\gamma_{3}:=\tfrac{\epsilon/(3c_{2})}{2L_{\phi}(1+2\tau_{\text{per}})\exp(L_{\phi}\theta^{\star})}$,$\gamma_{4}:=\tfrac{\epsilon/(3c_{2})}{2L_{\phi}\tau_{\text{per}}}$, $\gamma_{1}:=\min\{\gamma_{0},\gamma_{2},\gamma_{3},\gamma_{4},1\}$. Pick $\gamma\in(0,\gamma_{1})$ such that $t^{\star}:=\frac{\theta^{\star}}{\gamma}\in{\mathbb{N}}$. This can be done without loss of generality since $\theta^{\star}$ is a design parameter. Then, we bound (38) as
$\|\xi^{t}-\xi_{\text{avg}}^{t}\|\leq\frac{\epsilon}{c_{2}},$ (40)
for all $t\in\{0,\dots,t^{\star}\}$. As a consequence, since $\xi_{\text{avg}}^{t}\in\mathcal{B}_{r_{\xi}^{\prime}}$ for all $t\in\mathbb{N}$, it holds ${\xi}^{t}\in\mathcal{B}_{r_{\xi}^{\prime}+\epsilon}$ for all $t\in\{0,\dots,t^{\star}\}$, i.e., we have verified that the bounds (34) can be used into the interval $\{0,\dots,t^{\star}\}$. Moreover, the exponential law (31) and the expression of $\theta^{\star}$ (cf. (39)) ensure that it holds
@@
where in $(a)$ we combined (40) and (41). The inequality (42) guarantees that $\xi^{t^{\star}}\in\mathcal{B}_{\frac{\bar{\rho}}{2c_{2}}}$, hence we proved that the trajectories of (27) enters into $\mathcal{B}_{\frac{\bar{\rho}}{2c_{2}}}$ with linear rate. Next, in order to show that $\xi^{t}\in\mathcal{B}_{\bar{\rho}/2}$ for all $t\geq t^{\star}$, we divide the set of natural numbers in intervals as =$\mathbb{N}=\{0,\dots,t^{\star}\}\cup\{t^{\star},\dots,2t^{\star}\}\cup\dots$. Define $\psi_{\text{avg}}(k+t^{\star},\xi^{t^{\star}})$ as the solution to (27) for $\xi_{\text{avg}}^{0}=\xi^{t^{\star}}$ and $k\in\{0,\dots,t^{\star}\}$. Thus, at the beginning of each interval $\{t^{\star},\dots,2t^{\star}\}$, the initial condition of (27) coincides with the one of $\psi_{\text{avg}}(k+t^{\star},\xi^{t^{\star}})$ and lies into $\mathcal{B}_{\bar{\rho}/2}\subseteq\mathcal{B}_{r_{\xi}}$. Thus, we apply the same arguments above to guarantee that
$\|\xi^{k+t^{\star}}-\psi_{\text{avg}}(k+t^{\star},\xi^{t^{\star}})\|$ $\leq\epsilon$
$\psi_{\text{avg}}(2t^{\star},\xi^{t^{\star}})$ $\in\mathcal{B}_{(\bar{\rho}/2-\epsilon)/c_{2}},$
for all $\gamma\in(0,\gamma^{\star})$ and $k\in\{0,\dots,t^{\star}\}$. By using Lemma 3, we guarantee that system (27) cannot escape from $\mathcal{B}_{\bar{\rho}/2-\epsilon}$, namely $\xi_{\text{avg}}^{t}\in\mathcal{B}_{\bar{\rho}/2-\epsilon}$ for all $t\geq t^{\star}$. Thus, we get $\xi^{t}\in\mathcal{B}_{\bar{\rho}/2}$ for all $t\in\{t^{\star},\dots,2t^{\star}\}$. By recursively applying the same arguments above for each interval $\{jt^{\star},\dots,(j+1)t^{\star}\}$ with $j=2,3,\dots$ and using $\left\|x_{i}^{t}-x^{\star}\right\|\leq\left\|\xi^{t}\right\|$ for all $i\in\{1,\ldots,N\}$ and $t\in\mathbb{N}$, we get
@@
for all $i\in\{1,\ldots,N\}$ and $t\in\mathbb{N}$. The change of coordinates (6), (43), and the triangle inequality lead to
$\left\|w_{i}^{t}-x^{\star}\right\|$ $\leq\frac{\bar{\rho}}{2}+\delta\left\|d_{i}^{t}\right\|\stackrel{{\scriptstyle(a)}}{{\leq}}\frac{\bar{\rho}}{2}+\delta\sqrt{n},$ (44)
@@ V Numerical Computations on Distributed personalized optimization
In several engineering applications, a problem of interest consists of optimizing a performance metric while keeping into account user discomfort terms [9, 44]. In these scenarios, the user discomfort term is usually not known in advance but can be only accessed by measurements. Specifically, we associate to each agent $i\in\{1,\ldots,N\}$ a cost function in the form $f_{i}(w)=w^{\top}Q_{i}w+r_{i}^{\top}w+\log(\sum_{\ell=1}^{n}a_{i\ell}e^{b_{i\ell}w_{\ell}})$ with $Q_{i}=Q_{i}^{\top}\in^{n\times n}$, $r_{i}\in^{n}$ and $a_{i\ell},b_{i\ell}>0$, for all $\ell\in\{1,\ldots,n\}$. In the following, we provide different sets of simulations to study in detail the features of Extremum Seeking Tracking. In particular, each set consists of Monte Carlo simulations over $20$ randomly generated scenarios in which each one differs from the other in terms of cost functions, communication graphs, and algorithmic variables’ initialization. In particular, unless differently stated, in each trial the agents communicate according to Erdős-Rényi random graphs (see, e.g., [45]) with edge probabilities equal to $0.2$. As for the problem parameters, for each trial and all $i\in\{1,\ldots,N\}$, we generate each matrix $Q_{i}$ by pre- and post-multiplying a diagonal matrix (whose diagonal elements are randomly extracted from the interval $[10^{-3},5\cdot 10^{-3}]$ with uniform probability) with an orthonormal matrix (randomly generated by extracting its elements from the interval $[0,1]$ with uniform probability) and its transpose, respectively. Further, for all $i\in\{1,\ldots,N\}$ and $\ell\in\{1,\ldots,n\}$, we randomly extract the components of $r_{i}$ within the interval $[-10^{-2},3\cdot 10^{-2}]$ and the parameters $a_{i\ell},b_{i\ell}$ within the interval $[0,10^{-3}]$ with a uniform probability. In each set of simulations, we choose the parameters $\tau_{i_{p}}$ and $\phi_{i_{p}}$ according to the following procedure. For all $i\in\{1,\ldots,N\}$, we take $\phi_{i_{p}}=\dfrac{\pi}{4}\left(1+(-1)^{p}\right)$ for all $p=1,\dots,n$, while $\tau_{i_{p}}$ have been chosen as the first $\lfloor(n+1)/2\rfloor$ elements of odd numbers greater than $3$ since it is possible to show that such a set of frequencies satisfy (3). Roughly speaking, $\phi_{i,p}=0$ and $\phi_{i,p}=\pi/2$ are used to create orthogonal functions with the same frequency. It is important to note that this selection is not unique. As for the algorithm parameters $\gamma$ and $\delta$, due to the complexity of the dependencies of the bounds $\gamma^{\star}$ and $\delta^{\star}$ provided in Theorem 1, we choose them via a trial-and-error procedure. In each set of simulations, the performance is evaluated by providing graphical results involving the relative errors ${|\sum_{i}f_{i}(\bar{x}^{t})-f(x^{\star})|}/{|f(x^{\star})|}$ and ${\|\bar{x}^{t}-x^{\star}\|}/{\|x^{\star}\|}$ achieved along the trials of the Monte Carlo simulations, where $\bar{x}^{t}:=\frac{1}{N}\sum_{i=1}^{N}x_{i}^{t}$. Simulations are performed using DISROPT [46], a Python package based on MPI to encode and simulate distributed optimization algorithms.
@@ V-A Monte Carlo simulations for large-scale problems
First, we perform numerical simulations over large-scale problems with networks made of $N=250$ agents. We consider different optimization variable sizes, namely $n=10,20$. Part of these simulations has been run on the Marconi100 HPC Cluster of the Italian Cineca. We used $10$ nodes of the cluster and, for each node, we used $25$ cores and $4$ GPUs. The code has been adapted in order to perform part of the computation directly on GPUs. The results are shown in Fig. 2. In detail, as the number of agents increases, the Lipschitz constant of the system to be averaged increases too. Moreover, a larger domain of initial conditions also implies a potentially larger $L_{\phi}$ constant (cf. (33)). This implies smaller $\gamma^{\star}$, which, fixed the other parameters, makes the convergence slower. The decision variable dimension instead impacts the selection of the dither signal. A larger number of states implies a larger number of frequencies. This, in turn, means a longer time to estimate the gradient (cf. Lemma 1). Notice that, however, the accuracy of the final estimate is guaranteed by design. Indeed, since $\delta$ and $\gamma$ are designed on $\bar{\rho}$, the trajectories of (5) converge to a ball of radius $\bar{\rho}$ independently of the problem size.